OpenStudy (anonymous):

find an equation of the line passing through the given points. use function notation to write the equation. (-4,2) and (-6,-1)

OpenStudy (agent47):

equation of the line: y-y1=M(x-x1)

OpenStudy (anonymous):

can you solve the equation

OpenStudy (agent47):

M=(y2-y1)/(x2-x1) M=(-1-2)/(-6-(-4))=-3/-2=3/2

OpenStudy (agent47):

y-2=(3/2)(x-(-4)) y=(3/2)(x+4)+2 y=(3/2)x+6+2 y=(3/2)x+8

OpenStudy (anonymous):

y=3/2x+8 is the answer